Combined savings. Wanda makes $6000 more per year
than her husband does. Wanda saves 10% of her income
for retirement, and her husband saves 6%. If together they
save $5400 per year, then how much does each of them
make per year?

Answer by mananth(16949)

Combined savings. Wanda makes $6000 more per year
than her husband does. Wanda saves 10% of her income
for retirement, and her husband saves 6%. If together they
save $5400 per year, then how much does each of them
make per year?
Husband income = $x
Wanda income = x+6000
...
Saving
Husband =>0.06x
Wanda = >0.1(x+6000)
...
Total saving = 5400
...
0.06x+0.1(x+6000)=5400
0.06x+0.1x+600=5400
0.16x=5400-600
0.16x=4800
/0.16
x=4800/0.16
x=$30,000 ===========> Husband's income
x+6000=$36,000=======> Wanda's income